# Transverse Instability of Stokes Waves at Finite Depth

### v1

@@ 1. We claim that the mapping
$\tfrac{h_{\varepsilon}}{h}\mathcal{G}_{\varepsilon,\beta}=\partial_{z}\Psi(\varepsilon,\beta)|_{z=0}=\sum_{j=0}^{\infty}\sum_{\ell=0}^{\infty}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\partial_{z}\Psi_{j,\ell}|_{z=0}\quad\text{in }H^{s-1}(\mathbb{T}).$
Thus the mapping (2.35) is analytic. Because the trace theorem fails for $s\leq 1$, we shall provide a different argument for $s=\frac{1}{2}$ and then conclude the proof by interpolation.
Now we fix $s=\frac{1}{2}$. Then the coefficient $\Psi_{j,\ell}$ belongs to $H^{1}(\Omega_{0})$ and satisfies the bound (2.67). First, since $\Psi_{0,0}=\Theta|_{(\varepsilon,\beta)=(0,\beta^{0})}$, in view of the explicit formula (2.43), we have $\partial_{z}\Psi_{0,0}|_{z=0}\in H^{-\frac{1}{2}}(\mathbb{T})$ and $\|\partial_{z}\Psi_{0,0}|_{z=0}\|_{H^{-\frac{1}{2}}}\leq C\|f\|_{H^{\frac{1}{2}}}$. Next, for general $j$ and $k$ we will appeal to Lemma A.2 to make sense of $\partial_{z}\Psi_{j,k}|_{z=0}$. For brevity, we define
$w_{j,\ell}:=\sum_{k=0}^{j}\nu_{j-k}\Psi_{k,\ell}.$
In case $\ell=0$ and $j\geq 1$, we add $\Delta_{x,z}\sum_{k=0}^{j-1}\nu_{j-k}\Psi_{k,0}$ to both sides of (2.58) and use $\nu_{0}=\chi_{0}=1$ to obtain
$\Delta_{x,z}w_{j,0}=\beta^{0}\sum_{k=0}^{j}\chi_{j-k}\Psi_{k,0}+\partial_{x}^{2}\sum_{k=0}^{j-1}\nu_{j-k}\Psi_{k,0}\quad\text{in~{}}\Omega_{0}.$ (2.72)
More generally, for $\ell\geq 1$ we add $\Delta_{x,z}\sum_{k=0}^{j-1}\nu_{j-k}\Psi_{k,\ell}$ to both sides of (2.59) to obtain
$\Delta_{x,z}w_{j,\ell}=\sum_{k=0}^{j}\chi_{j-k}\Psi_{k,\ell-1}+\beta^{0}\sum_{k=0}^{j}\chi_{j-k}\Psi_{k,\ell}+\partial_{x}^{2}\sum_{k=0}^{j-1}\nu_{j-k}\Psi_{k,\ell}\quad\text{in~{}}\Omega_{0}.$ (2.73)
Hence, applying the estimate (A.7) in Lemma A.2 to $u=\nabla w_{j,\ell}$, we have $\tilde{\gamma}(\nabla w_{j,\ell})\in(H^{-\frac{1}{2}}(\mathbb{T}))^{2}$ and the estimate
$\|\tilde{\gamma}(\nabla w_{j,\ell})\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}$ $\leq C\|\nabla w_{j,\ell}\|_{L^{2}(\Omega_{0})}+C\left\|\sum_{k=0}^{j}\chi_{j-k}\Psi_{k,\ell-1}+\beta^{0}\sum_{k=0}^{j}\chi_{j-k}\Psi_{k,\ell}\right\|_{L^{2}(\Omega_{0})}$
$\qquad+C\left\|\partial_{x}\sum_{k=0}^{j-1}\nu_{j-k}\Psi_{k,\ell}\right\|_{L^{2}(\Omega_{0})}$
with the convention that $\Psi_{j,-1}=0$, where $C$ depends only on $h$. Then, using the definition of $w_{j,\ell}$ and invoking the estimates (2.67) and (2.60) for $\|\Psi_{j,\ell}\|_{H^{1}(\Omega_{0})}$, as well as the bounds (2.55) for $\chi_{j-k}$ and $\nu_{j-k}$, we find
$\|\tilde{\gamma}(\nabla w_{j,\ell})\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}\leq\frac{C_{1}}{A-1}\|f\|_{H^{\frac{1}{2}}(\mathbb{T})}(AM)^{j+\ell},\quad j,~{}\ell\geq 0,$ (2.74)
recalling that $A>1$ and $M>1$, where $C_{1}$ depends only $(\beta^{0},h)$.
From (2.60) we have $\|\Psi_{0,0}\|_{H^{1}(\Omega_{0})}\leq M_{0}\|f\|_{H^{\frac{1}{2}}(\mathbb{T})}$, $M_{0}=M_{0}(\beta^{0},h)$. Then, the first equation in (2.57) implies $\Delta\Psi_{0,0}\in H^{1}(\Omega_{0})\subset L^{2}(\Omega_{0})$, so that $\nabla\Psi_{0,0}\in H_{\operatorname{div}}(\Omega_{0})$. Therefore, we can apply Proposition A.1 and Lemma A.2 to have $\tilde{\gamma}(\nabla\Psi_{0,0})={\gamma}(\nabla\Psi_{0,0})$ and $\|\tilde{\gamma}(\nabla\Psi_{0,0})\|_{H^{-\frac{1}{2}}(\mathbb{T})}\leq C_{2}\|f\|_{H^{\frac{1}{2}}(\mathbb{T})}$ for some constant $C_{2}=C_{2}(\beta^{0},h)\geq C_{1}$.
Now we prove by induction on $j+\ell$ that
$\|\tilde{\gamma}(\nabla\Psi_{j,\ell})\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}\leq C_{2}\|f\|_{H^{\frac{1}{2}}}(AM)^{j+\ell},\quad j,~{}\ell\geq 0.$ (2.75)
Indeed, the case $j+\ell=0$ is true. Suppose by induction that (2.75) holds for all pairs $(j,\ell)$ with $j+\ell\leq L-1$, and consider $j+\ell=L$. By the definition of $w_{j,\ell}$ we have $\Psi_{j,\ell}=w_{j,\ell}-\sum_{k=0}^{j-1}\nu_{j-k}\Psi_{k,\ell}$, so that (2.74) and the induction hypothesis yield
$\|\tilde{\gamma}(\nabla\Psi_{j,\ell})\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}$ $\leq\|\tilde{\gamma}(\nabla w_{j,\ell})\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}+\sum_{k=0}^{j-1}|\nu_{j-k}\|\tilde{\gamma}(\nabla\Psi_{k,\ell})\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}$
$\leq\frac{C_{2}}{A-1}\|f\|_{H^{\frac{1}{2}}(\mathbb{T})}(AM)^{j+\ell}+\sum_{k=0}^{j-1}M^{j-k+1}C_{2}\|f\|_{H^{\frac{1}{2}}}(AM)^{k+\ell}$
$\leq\frac{C_{2}}{A-1}\|f\|_{H^{\frac{1}{2}}(\mathbb{T})}(AM)^{j+\ell}+\frac{C_{2}M}{A-1}\|f\|_{H^{\frac{1}{2}}}(AM)^{j+\ell}$
$\leq C_{2}\|f\|_{H^{\frac{1}{2}}}(AM)^{j+\ell},$
where in the last inequality we have used the fact from (2.66) that $A-1>1+M$. The proof of (2.75) is complete.
Combining (2.56) with (2.75), we deduce that
the series $\text{the series ~{}}\sum_{j=0}^{\infty}\sum_{\ell=0}^{\infty}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\tilde{\gamma}(\nabla\Psi_{j,\ell})\text{~{}converges to some~{}}g~{}\text{in~{}}(H^{-\frac{1}{2}}(\mathbb{T}))^{2}$ (2.76)
provided $|\varepsilon|<(AM)^{-1}$ and $|\beta-\beta^{0}|<(AM)^{-1}$. We will show that $g=\tilde{\gamma}(\nabla\Psi(\varepsilon,\beta))$. To that end we define
$u_{m,n}:=\begin{bmatrix}\partial_{x}\Psi^{(m,n)}\\ H^{(m)}\partial_{z}\Psi^{(m,n)}\end{bmatrix}=H^{(m)}\nabla\Psi^{(m,n)}+\begin{bmatrix}(1-H^{(m)})\partial_{x}\Psi^{(m,n)}\\ 0\end{bmatrix},$ (2.77)
where $H^{(m)}$ is given in (2.68). Combining Lemma A.2 (i)(c) with (2.76), we obtain
[…]
On the other hand, by the first equations in (2.57), (2.58) and (2.59), we find
$\operatorname{div}u_{m,n}=\partial_{x}^{2}\Psi^{(m,n)}+H^{(m)}\partial_{z}^{2}\Psi^{(m,n)}=\beta^{0}\chi^{(m)}\Psi^{(m,n)}+(\beta-\beta^{0})\chi^{(m)}\Psi^{(m,n-1)}\in H^{1}(\Omega_{0}),$
which can be proven similarly to (2.69). In addition,
$u:=\begin{bmatrix}\partial_{x}\Psi(\varepsilon,\beta)\\ \big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2}\partial_{z}\Psi(\varepsilon,\beta)\end{bmatrix}$
satisfies $\operatorname{div}u=\beta\tilde{J}\Psi(\varepsilon,\beta)\in H^{1}(\Omega_{0})$ in light of (2.44). It follows that $u_{m,n},\ u\in H_{\operatorname{div}}(\Omega_{0})$, and hence (A.2) implies
$\|\gamma(u_{m,n}-u)\|_{(H^{-\frac{1}{2}}(\mathbb{T}))^{2}}\leq C\|u_{m,n}-u\|_{L^{2}(\Omega_{0})}+C\|\operatorname{div}(u_{m,n}-u)\|_{L^{2}(\Omega_{0})}\to 0\quad\text{as~{}}m,n\to\infty,$
where we have used the fact from (2.56) that $\Psi^{(m,n)}\to\Psi(\varepsilon,\beta)$ in $H^{1}(\Omega_{0})$. Recalling our definition (A.6) of $\tilde{\gamma}$ on $H_{\operatorname{div}}+\tilde{H}_{\operatorname{div}}$, we obtain
$\lim_{m,n\to\infty}\tilde{\gamma}(u_{m,n})=\tilde{\gamma}(u)=\big{(}\frac{h}{h_{\varepsilon}}\big{)}^{2}\tilde{\gamma}(\nabla\Psi(\varepsilon,\beta))\quad\text{in~{}}(H^{-\frac{1}{2}}(\mathbb{T}))^{2},$
where the second equality follows from a decomposition analogous to (2.77).
In conjunction with (2.78) this yields $g=\tilde{\gamma}(\nabla\Psi(\varepsilon,\beta))$ as claimed. It means that
$\tilde{\gamma}(\nabla\Psi(\varepsilon,\beta))=\sum_{j=0}^{\infty}\sum_{\ell=0}^{\infty}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\tilde{\gamma}(\nabla\Psi_{j,\ell})\quad\text{in~{}}(H^{-\frac{1}{2}}(\mathbb{T}))^{2}.$ (2.79)
Due to (2.45), this yields the analyticity of the mapping (2.35) when $s=\frac{1}{2}$.
Finally, in order to generalize to $s\geq\frac{1}{2}$, we shall interpolate. We recall (2.71) and view $M_{0}$, $A$ and $M$ as functions of $s\geq\frac{1}{2}$. We fix any $s_{0}>1$ and use the standard trace theorem to obtain
$\|\gamma(\nabla\Psi_{j,\ell})\|_{(H^{s_{0}-1}(\mathbb{T}))^{2}}\leq N(s_{0})\|\Psi_{j,\ell}\|_{H^{s_{0}+\frac{1}{2}}(\Omega_{0})}\leq N(s_{0})M_{0}(s_{0})\big{(}A(s_{0})M(s_{0})\big{)}^{j+\ell}\|f\|_{H^{s_{0}}(\mathbb{T})}.$
Thus we can view $\gamma(\nabla\Psi_{j,\ell})=\tilde{\gamma}(\nabla\Psi_{j,\ell})$ as an operator in $\mathcal{L}(H^{s_{0}}(\mathbb{T}),H^{s_{0}-1}(\mathbb{T}))^{2})$. In light of this inequality and (2.75), we obtain by interpolation that
$\|\tilde{\gamma}(\nabla\Psi_{j,\ell})\|_{\mathcal{L}(H^{s}(\mathbb{T}),(H^{s-1}(\mathbb{T}))^{2})}\leq C_{2}^{\kappa}\big{(}A(\tfrac{1}{2})M(\tfrac{1}{2})\big{)}^{\kappa(j+\ell)}\big{(}N_{0}(s_{0})M_{0}(s_{0})\big{)}^{1-\kappa}\big{(}A(s_{0})M(s_{0})\big{)}^{(1-\kappa)(j+\ell)}$
for any $s=\kappa\frac{1}{2}+(1-\kappa)s_{0}\in(\frac{1}{2},s_{0})$, $\kappa\in(0,1)$. Consequently the series $\sum_{j=0}^{\infty}\sum_{\ell=0}^{\infty}\varepsilon^{j}(\beta-\beta^{0})^{\ell}\tilde{\gamma}(\nabla\Psi_{j,\ell})$ converges in $(H^{s-1}(\mathbb{T}))^{2})$ provided $|\varepsilon|$ and $|\beta-\beta^{0}|$ are small enough. The sum of this series must equal $\tilde{\gamma}(\nabla\Psi(\varepsilon,\beta))$ thanks to (2.79). Therefore, the mapping (2.35) is analytic in a neighborhood of $\varepsilon=0$ and $\beta=\beta^{0}$ for any $s\geq\frac{1}{2}$. The proof of Theorem 2.6 is now complete.
@@ 2.4. The reformulated instability problem
The proof of Proposition 2.10 is given in Appendix A.2 of [7], noting that their $f|_{\varepsilon}$, $\mathfrak{p}_{\varepsilon}$, $p_{\varepsilon}$, and $a_{\varepsilon}$ correspond to $(h_{\varepsilon}-h)$, $(\zeta-x)$, $(p-c_{0})$, and $[(1+q)/\zeta^{\prime}-1]$, respectively, in this work. ∎
@@ 3.1. Resonance condition
$-c_{0}(m+1)+((m+1)^{2}+\beta)^{\frac{1}{4}}\tanh^{\frac{1}{2}}\left(h\left((m+1)^{2}+\beta\right)^{\frac{1}{2}}\right)=c_{0}m-(m^{2}+\beta)^{\frac{1}{4}}\tanh^{\frac{1}{2}}\left(h\left(m^{2}+\beta\right)^{\frac{1}{2}}\right).$ (3.4)
For our purposes, we choose the simplest resonance $m=1$. Then, we have the following proposition on the existence, uniqueness and asymptotics of the resonant value of $\beta$.
@@
where the first inequality follows from the monotonicity of $\tanh^{\frac{1}{2}}(\xi)$ for $\xi>0$. Consequently, for each $h>0$, there exists $30<\beta_{*}<$ 3 such that $F(\beta_{*},h)=0$. Because $\partial_{\beta}F<0$, we deduce that $\beta_{*}(h)$ is the unique positive zero of $F(\cdot,h)$ This proves (i). By the Analytic Implicit Function Theorem, $\beta_{*}(h)$ is real analytic for $h>0$, proving (ii).
@@
and upon squaring both sides, while using the bound $30<\beta_{*}(h)<$ 3, we find
@@
From the monotonicity of $\tanh$, we have $\tanh(h)/\tanh(wh)<1$ for $h>0,\ w>1$. Thus
@@
For $30<\beta<$ 3, we have the bounds
@@
Now $01-4^{\frac{3}{4}}/3>$ 0 and $\textrm{sech}(2h)/\textrm{sech}(h)\rightarrow 0$ as $h\rightarrow\infty$, so that there exists $H_{0}>0$ such that
@@
Consequently, $\partial_{h}F(\beta,h)>0$ for all $30<\beta<$ 3 and $h>H_{0}$. In particular, $(\partial_{h}F)(\beta_{*}(h),h)>0$ for $h>H_{0}$. However, $\partial_{\beta}F(\beta,h)<0$ for all $\beta>0$ and $h>0$, so that $\beta_{*}^{\prime}(h)=-\frac{(\partial_{h}F)(\beta_{*}(h),h)}{(\partial_{\beta}F)(\beta_{*}(h),h)}>0$ for $h>H_{0}$, proving our claim.
@@
$0\displaystyle 3-(1+\beta_{*,\infty})^{\frac{1}{4}}-(4+\beta_{*,\infty})^{\frac{1}{4}}=$ 0 . (3.7)
@@ 3.2. Kato’s perturbed basis
[…]
@@ 4.1. Third-order expansions of $\mathcal{L}_{\varepsilon,\beta}$
The coefficients $r_{1,1},\dots,p_{3,3}$ are given in Prop. 2.10. The nonlocal operators $R_{j,\beta}$ are Fourier multipliers. We explicitly compute them for $0\leq j\leq 3$ in the following proposition.
@@
where the coefficients are given in Prop. 2.10 and Prop. 2.2. A Taylor expansion leads to
@@
$&\Delta_{x,z}\Theta^{2}-\beta\Theta^{2}=\beta\biggr{[}\frac{C_{2(h+z)}}{2C_{h}^{2}}\zeta_{1,1}^{2}+\Big{(}4\frac{C_{2(h+z)}}{C_{2h}}\zeta_{2,2}+\frac{1}{2C_{h}^{2}}\zeta_{1,1}^{2}\Big{)}\cos(2x)\biggr{]}\Theta^{0}\\ &\hskip 88.2037pt+2\beta\frac{C_{h+z}}{C_{h}}\zeta_{1,1}\cos x\Theta^{1}\quad\text{in }\{(x,z):-h<z<0\},\\ &\Theta^{2}(x,0)=0,\\ &\partial_{z}\Theta^{2}(x,-h)=h_{2}\partial_{z}^{2}\Theta^{0}(x,-h),\end{aligned}\end{cases}$ (4.13)
$\Delta_{x,z}\Theta^{3}-\beta\Theta^{3}&=2\beta\biggr{[}\Big{(}\frac{d_{2}S_{z}}{C_{h}^{2}}\zeta_{1,1}+\frac{C_{3(h+z)}}{C_{h}C_{2h}}\zeta_{1,1}\zeta_{2,2}+\frac{C_{h+z}}{C_{h}}\zeta_{3,1}\Big{)}\cos x+\Big{(}\frac{C_{h+z}}{C_{h}C_{2h}}\zeta_{1,1}\zeta_{2,2}\\ &\hskip 24.18501pt+3\frac{C_{3(h+z)}}{C_{3h}}\zeta_{3,3}\Big{)}\cos(3x)\biggr{]}\Theta^{0}+\beta\biggr{[}\frac{C_{2(h+z)}}{2C_{h}^{2}}\zeta_{1,1}^{2}+\Big{(}4\frac{C_{2(h+z)}}{C_{2h}}\zeta_{2,2}\\ &\hskip 24.18501pt+\frac{1}{2C_{h}^{2}}\zeta_{1,1}^{2}\Big{)}\cos(2x)\biggr{]}\Theta^{1}+2\beta\frac{C_{h+z}}{C_{h}}\zeta_{1,1}\cos x\Theta^{2}\quad\text{in}\{(x,z):-h<z<0\},\end{aligned}\\ \Theta^{3}(x,0)=0,\\ \partial_{z}\Theta^{3}(x,-h)=h_{2}\partial_{z}^{2}\Theta^{1}(x,-h).\end{cases}$ (4.14)

### v2

@@ 2.3. The transformed Dirichlet-Neumann operator
$\partial_{z}\Psi(\varepsilon,\beta)|_{z=0}:(\varepsilon,\beta)\ni\mathcal{O}_{s}\to H^{s-1}(\mathbb{T})\quad\text{is~analytic}$ (2.58)
for some neighborhood $\mathcal{O}_{s}$ of $(\varepsilon,\beta)=(0,\beta_{0})$. To treat the case $s=\frac{1}{2}$ we consider the vector field
$u=\begin{bmatrix}\partial_{x}\Psi(\varepsilon,\beta)\\ \big(\frac{h}{h_{\varepsilon}}\big)^{2}\partial_{z}\Psi(\varepsilon,\beta)\end{bmatrix}$.
Because $\operatorname{div}u=\beta\tilde{J}\Psi(\varepsilon,\beta)\in H^{1}(\Omega_{0})\subset L^{2}(\Omega_{0})$ in view of (2.53), we have that $u\in H_{\operatorname{div}}(\Omega_{0})$, defined by (2.37). Therefore, by virtue of Proposition 2.7, the trace $\gamma(u)=:\big(\frac{h}{h_{\varepsilon}}\big)^{2}\partial_{z}\Psi(\varepsilon,\beta)|_{z=0}$ is well-defined in $H^{-\frac{1}{2}}(\mathbb{T})$. Therefore, (2.58) is still valid for $s=\frac{1}{2}$. Viewing $\Psi(\varepsilon,\beta)$ as the bounded linear operator $f\mapsto\Psi$ solution of (2.53), we deduce with the aid of linear interpolation that the map
$f\mapsto\partial_{z}\Psi(\varepsilon,\beta)|_{z=0}:(\varepsilon,\beta)\ni\mathcal{O}_{s}\to\mathcal{L}(H^{s}(\mathbb{T}),H^{s-\frac{1}{2}}(\mathbb{T}))$
is analytic for all $s\geq\frac{1}{2}$. This concludes the proof of the analyticity (2.35) of $\mathcal{G}_{\varepsilon,\beta}$. $\square$
@@ 2.4. The reformulated instability problem
[…]
@@ 3.1. Resonance condition
$c_{0}(m+1)+((m+1)^{2}+\beta)^{\frac{1}{4}}\tanh^{\frac{1}{2}}\left(h\left((m+1)^{2}+\beta\right)^{\frac{1}{2}}\right)=c_{0}m-(m^{2}+\beta)^{\frac{1}{4}}\tanh^{\frac{1}{2}}\left(h\left(m^{2}+\beta\right)^{\frac{1}{2}}\right).$ (3.4)
For our purposes, we choose the simplest resonance $m=1$, which leads to instabilities with the largest growth rates for small Stokes waves as suggested by numerical computations [34]. Then, we have the following proposition on the existence, uniqueness and asymptotics of the resonant value of $\beta$.
@@
where the first inequality follows from the monotonicity of $\tanh^{\frac{1}{2}}(\xi)$ for $\xi>0$. Consequently, for each $h>0$, there exists $0<\beta_{*}<3$ such that $F(\beta_{*},h)=0$. Because $\partial_{\beta}F<0$, we deduce that $\beta_{*}(h)$ is the unique positive zero of $F(\cdot,h)$ . This proves (i). By the Analytic Implicit Function Theorem, $\beta_{*}(h)$ is real analytic for $h>0$, proving (ii).
@@
and upon squaring both sides, while using the bound $0<\beta_{*}(h)<3$, we find
@@
From the monotonicity of $\tanh$, we have $\tanh(h)/\tanh(wh)<1$ for h>0,w>1h>0,\ w>1. Thus
@@
For $0<\beta<3$, we have the bounds
@@
Now $1-4^{\frac{3}{4}}/3>0$ and $\textrm{sech}(2h)/\textrm{sech}(h)\rightarrow 0$ as $h\rightarrow\infty$, so that there exists $H_{0}>0$ such that
@@
Consequently, $\partial_{h}F(\beta,h)>0$ for all $0<\beta<3$ and $h>H_{0}$. In particular, $(\partial_{h}F)(\beta_{*}(h),h)>0$ for $h>H_{0}$. However, $\partial_{\beta}F(\beta,h)<0$ for all $\beta>0$ and $h>0$, so that $\beta_{*}^{\prime}(h)=-\frac{(\partial_{h}F)(\beta_{*}(h),h)}{(\partial_{\beta}F)(\beta_{*}(h),h)}>0$ for $h>H_{0}$, proving our claim.
@@
$3-(1+\beta_{*,\infty})^{\frac{1}{4}}-(4+\beta_{*,\infty})^{\frac{1}{4}}=0.$ (3.7)
@@ 3.2. Kato’s perturbed basis
$P_{\varepsilon,\delta}=-\frac{1}{2\pi i}\int_{\Gamma}(\mathcal{L}_{\varepsilon,\beta_{*}+\delta}-\lambda)^{-1}d\lambda\in\mathcal{L}\big((L^{2}(\mathbb{T}))^{2},(H^{1}(\mathbb{T}))^{2}\big).$ (3.17)
@@ 4.1. Third-order expansions of $\mathcal{L}_{\varepsilon,\beta}$
The coefficients $r_{1,1},\dots,p_{3,3}$ are given in Prop. 2.9. The nonlocal operators $R_{j,\beta}$ are Fourier multipliers. We explicitly compute them for $0\leq j\leq 3$ in the following proposition.
@@
where the coefficients are given in Prop. 2.9 and Prop. 2.2. A Taylor expansion leads to
@@
$\begin{cases}\begin{aligned} &\Delta_{x,z}\Theta^{2}-\beta\Theta^{2}=\beta\biggr[\frac{C_{2(h+z)}}{2C_{h}^{2}}\zeta_{1,1}^{2}+\Big(4\frac{C_{2(h+z)}}{C_{2h}}\zeta_{2,2}+\frac{1}{2C_{h}^{2}}\zeta_{1,1}^{2}\Big)\cos(2x)\biggr]\Theta^{0}\\ &\hskip 88.2037pt+2\beta\frac{C_{h+z}}{C_{h}}\zeta_{1,1}\cos x\Theta^{1}\quad\text{in }\{(x,z):-h<z<0\},\\ &\Theta^{2}(x,0)=0,\\ &\partial_{z}\Theta^{2}(x,-h)=h_{2}\partial_{z}^{2}\Theta^{0}(x,-h),\end{aligned}\end{cases}$ (4.13)
$\begin{cases}\begin{aligned} \Delta_{x,z}\Theta^{3}-\beta\Theta^{3}&=2\beta\biggr[\Big(\frac{d_{2}S_{z}}{C_{h}^{2}}\zeta_{1,1}+\frac{C_{3(h+z)}}{C_{h}C_{2h}}\zeta_{1,1}\zeta_{2,2}+\frac{C_{h+z}}{C_{h}}\zeta_{3,1}\Big)\cos x+\Big(\frac{C_{h+z}}{C_{h}C_{2h}}\zeta_{1,1}\zeta_{2,2}\\ &\hskip 24.18501pt+3\frac{C_{3(h+z)}}{C_{3h}}\zeta_{3,3}\Big)\cos(3x)\biggr]\Theta^{0}+\beta\biggr[\frac{C_{2(h+z)}}{2C_{h}^{2}}\zeta_{1,1}^{2}+\Big(4\frac{C_{2(h+z)}}{C_{2h}}\zeta_{2,2}\\ &\hskip 24.18501pt+\frac{1}{2C_{h}^{2}}\zeta_{1,1}^{2}\Big)\cos(2x)\biggr]\Theta^{1}+2\beta\frac{C_{h+z}}{C_{h}}\zeta_{1,1}\cos x\Theta^{2}\quad\text{in }\{(x,z):-h<z<0\},\end{aligned}\\ \Theta^{3}(x,0)=0,\\ \partial_{z}\Theta^{3}(x,-h)=h_{2}\partial_{z}^{2}\Theta^{1}(x,-h).\end{cases}$ (4.14)